OpenStudy (anonymous):

PLEASE HELP! Solve 5x + 2y = 7 10x + 4y = 14 through Elimination

OpenStudy (anonymous):

5x+2y = 7 ..........(1) 10x + 4y = 14.........(2) (1) x 2 10x + 4y = 14........(3) (3)-(2) that gives 0 what a weird question

OpenStudy (anonymous):

Does it have infinite solutions?

OpenStudy (anonymous):

they are in fact the same equation so any value of x and y works

OpenStudy (anonymous):

oh ohk so would it be Infinite Solutions?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you ! c:

OpenStudy (phi):

yes, both equations are really the same equation. If you find a solution for one of them, you have a solution for the other because each equation is a "line", there are lots of pairs of (x,y) that solve the equation (all the points on the line)

OpenStudy (anonymous):

thanks ! I had trouble with this one but I got it now thank you guys c: